# Don’t look back: Ordering and defect cloaking in non-reciprocal lattice XY models

### v1

@@ Abstract
We present a detailed analytical and numerical examination, on square and triangular lattices, of the non-reciprocal planar spin model introduced in Dadhichi et al., Phys. Rev. E 101, 052601 (2020). We show that in principle the effect of lattice anisotropy should persist at large scales, leading to a “mass” for the angle field of the spins, and behaviour not in the “Malthusian Toner-Tu” universality class. Numerically, however, we find power-law scaling of long-wavelength equal-time correlators in the polar-ordered phase of our lattice model. The mass, if present, is very small. Focussing on topological defects, we show numerically that defect interactions are highly anisotropic with respect to the mean ordering direction. In particular, the constituents of a ±$\pm 1$ pair are shielded from each other in a class of configurations, deferring their annihilation and allowing time for the nucleation of further defects. The result, we show numerically, is the destruction of the polarised phase via an aster apocalypse reminiscent of that found by Besse et al., Phys. Rev. Lett. 129, 268003 (2022), for the Malthusian Toner-Tu equation.
@@ III The ordered state: Numerics
Figure 2: Two-point correlation function $C(r)$ for various system sizes $L$, with distance $r$ in units of $L/2\pi$, for (a) square and (b) triangular lattice, approaches a constant nonzero value for large $r$, indicating long-range order. Inset in each case shows magnetization $m(L)\equiv\sqrt{C(r=L)}$ as a function of $L^{-1}$. Measured structure factors $S_{\perp}$ as functions of $k_{\parallel}$ and $k_{\perp}$ for $L=512$ for (c) square and (d) triangular lattice show small-$k$ dependences consistent with $k_{\perp}^{-6/5}$ (and $\sim k_{\parallel}^{-2}$ for all but the smallest $k_{\parallel}$). The analysis is performed at reduced temperature $T=0.2$ and nonreciprocity strength $\mathcal{A}=1.5$, with time averaging over five independent noise realizations.
@@
We first examine the orientationally ordered state of our model. Initialising our system in an ordered state, we measure isotropically averaged spatial spin correlations $C(r)=\langle{\bf p}(r)\cdot{\bf p}(0)\rangle_{t}$ the angular brackets with a subscript denotes an average over that variable, for different system sizes $L$, nonreciprocity $\mathcal{A}$, and noise strength $T$. We benchmark our code by measuring these quantities for $\mathcal{A}=0$ which, as expected, displays quasi-long-range order below a defect-unbinding transition at temperature $T_{KT}\simeq 0.8$ for a $128\times 128$ square lattice.. For $\mathcal{A}\neq 0$, we find that $C(r)$ has a non-zero value at large $r$ signalling the existence of long-range order. That, in itself, is not too surprising: as we will discuss in the next section, when the order parameter space and the real space coincide, as it does in our model for $\mathcal{A}\neq 0$, lattice anisotropy explicitly breaks the continuous rotation symmetry of the model to a discrete one, even in equilibrium. Therefore, at low temperatures, we would expect long-range order in any such model. However, the reduction to a discrete symmetry would imply that the static structure factor of angular fluctuations $S_{\perp}({\bf k})\equiv\langle|\hat{\mathbf{p}}_{\perp}({\bf k},t)|^{2}\rangle$, where $\hat{\mathbf{p}}_{\perp}$ is the component of $\hat{\mathbf{p}}$ perpendicular to the instantaneous macroscopic ordering direction and ${\bf k}$ is the wavevector, should go to a $|{\bf k}|-$independent constant at small $|{\bf k}|$. We plot $S_{\perp}(k_{\parallel})$ and $S_{\perp}(k_{\perp})$, where $k_{\parallel}$ and $k_{\perp}$ are the components of ${\bf k}$ along and orthogonal to the ordering direction respectively, in Fig. 2c-d. While we find a clear departure from the $1/k_{\perp}^{2}$ expected when $\mathcal{A}=0$, they show a dependence on wavenumber consistent with a power law at least in the regime we have explored. The “mass” (see the next section) presumably conferred on the transverse fluctuations by the identification of the order parameter and lattice directions is not detected at these scales. In fact, Fig. 2c-d is consistent with the findings of [20].
@@ IV From lattice to continuum: the lingering effects of lattice anisotropy
where we have used the fact that the rank $n$ tensor $\sum_{a}[{\bf e}^{a}]^{\otimes n}$ vanishes for odd $n$. Writing $\hat{\mathbf{p}}({\bf x},t)\equiv\{\cos\theta({\bf x},t),\sin\theta({\bf x},t)\}$, where we have defined an angle field $\theta({\bf x},t)$ whose values at lattice sites $i$ correspond to $\theta^{i}$, this implies that on a square lattice,
@@
Additionally, reflection symmetry ensures that it is invariant under $y\to-y$, $\theta\to-\theta$ [25]. While (10) is symmetric under these transformations for terms up to $\mathcal{O}(\sigma)$, it is clearly not symmetric under (11) for terms at higher orders in $\sigma$. For instance, the term $\propto\sigma^{3}/6$ clearly breaks this symmetry at $\mathcal{O}(\nabla)$ in a way that modifies the scaling behaviour qualitatively (a hyperscaling relation that relates two relevant nonlinearities of the flocking model in isotropic space is no longer valid). Therefore, the microscopic model proposed in [17] cannot describe the spin wave dynamics of flocks in a fully rotation-invariant space. Second, one would expect the coupling of the order parameter to lattice directions to lead to an aligning field that favours alignment along any one of those directions, and thus a restoring torque proportional to the deviation $\theta$ of the orientation from that direction [19]. Such a term is apparently absent in the dynamics as derived above, but fluctuations may generate it. Though the interaction in [17] doesn’t have a special aligning direction, unlike traditional models with crystal fields [19, 18], the $\mathcal{O}(\sigma^{2})$ term, expanded to $\mathcal{O}(\theta^{3})$, is $\partial_{x}^{2}\theta+\theta^{2}(\partial_{y}^{2}\theta-\partial_{x}^{2}\theta)$. The first term makes the diffusivity anisotropic and we now argue that the second one may yield a contribution $\propto\theta$, whose presence would imply a finite relaxation time for even infinite wavelength angular fluctuations (if the sign of this term is negative). To see this notice that, within a self-consistent Gaussian decoupling, one can replace pairs of fields by their averages and write $\theta^{2}(\partial_{y}^{2}\theta-\partial_{x}^{2}\theta)\approx\langle\theta^{2}\rangle(\partial_{y}^{2}-\partial_{x}^{2})\theta+2\theta(\langle\theta\partial_{y}^{2}\theta\rangle-\langle\theta\partial_{x}^{2}\theta\rangle)$. The term within parentheses does not generically vanish in a state with long-range order, for which the correlators $\langle\theta\partial_{y}^{2}\theta\rangle$ and $\langle\theta\partial_{x}^{2}\theta\rangle$ are in general unequal: the “anisotropy exponent” $\zeta$, has no reason to remain $1$ and does not do so even in the Toner-Tu theory. This places the models considered in [17, 15, 16] within the active clock model universality class. It was conclusively shown in [18] that, unlike passive clock models which at high enough lattice coordination numbers and at finite noise strengths behave as isotropic spin systems, active $q$-state clock models remain only $q$-fold symmetric, and are thus distinct from polar active systems in isotropic space, for any $q$ no matter how large. Since order in the presence of anchoring to lattice directions should be long-ranged, the quasi-long-range ordered state reported in [15] is likely to be a finite-size effect.
@@
Since the interaction in [17] corresponds to the one we consider here to $\mathcal{O}(\sigma)$, and since the terms we identified as breaking isotropy all arise at higher order in $\sigma$, one might imagine that the spin wave theory of our model belongs to the Malthusian Toner-Tu universality class. We now argue that this is not the case. To do so, we use (3) to extend (7) to $\mathcal{O}(\nabla^{4})$ on square lattices. The angle field equation corresponding to this is
$\partial_{t}\theta=2\Bigg{(}\mathcal{A}(\cos\theta\partial_{x}\theta+\sin\theta\partial_{y}\theta)+\frac{J}{2}\nabla^{2}\theta+\frac{\mathcal{A}}{6}\{\cos\theta[\partial_{x}^{3}\theta-(\partial_{x}\theta)^{3}]+\sin\theta[\partial_{y}^{3}\theta-(\partial_{y}\theta)^{3}]\}\\ +\frac{J}{24}\{\partial_{x}^{4}\theta+\partial_{y}^{4}\theta-6[(\partial_{x}\theta)^{2}\partial_{y}^{2}\theta+(\partial_{y}\theta)^{2}\partial_{x}^{2}\theta]\}\Bigg{)}+\sqrt{2T}\xi_{\theta}\,.$ (12)
@@
The terms $\propto\mathcal{A}/6$ and $J/24$ in the large parentheses are not symmetric under (11), for distinct reasons. First, we discuss the term $\propto J/24$ which is symmetric under $\theta\to\theta+\theta_{0}$. This is because space and spin are decoupled in our model when $\mathcal{A}=0$, and spin rotation symmetry is preserved irrespective of the spatial anisotropy. The spatial gradient structure reflects the four-fold anisotropy of the lattice. The term $\propto\mathcal{A}/6$ is more interesting: the nonlinearities $\cos\theta(\partial_{x}\theta)^{3}$ and $\sin\theta(\partial_{y}\theta)^{3}$ have structures that can renormalise the coefficients of the terms $\cos\theta(\partial_{x}\theta)$ and $\sin\theta(\partial_{y}\theta)$ (as can be seen heuristically by replacing $(\partial_{x}\theta)^{2}$ and $(\partial_{y}\theta)^{2}$ by their averages, which yields $3\cos\theta\langle(\partial_{x}\theta)^{2}\rangle\partial_{x}\theta$ and $3\sin\theta\langle(\partial_{y}\theta)^{2}\rangle\partial_{y}\theta$). As argued earlier, $\langle(\partial_{x}\theta)^{2}\rangle$ and $\langle(\partial_{y}\theta)^{2}\rangle$ need not be equal in general (though they are equal within Gaussian fluctuations about our mean-field ordered state). This implies that the coefficient of the terms $\sin\theta\partial_{y}\theta$ and $\cos\theta\partial_{x}\theta$ need not remain equal in a renormalised theory, breaking the symmetry of our model under (11) even at $\mathcal{O}(\nabla)$. Since both of these terms contain relevant nonlinearities [25] (with a symmetry-mandated hyperscaling relation keeping the ratio of their coefficients fixed under renormalisation in flocks in isotropic space), the hydrodynamic behaviour of the ordered phase of our model cannot belong to the Toner-Tu universality class.
Strictly speaking, the important question of whether the large-scale, long-time physics of our model is the same as that in [17] remains unresolved. Gaussian decoupling arguments as applied above to the model of [17] do not yield a term $\propto\theta$ in our equation (12). As the dynamics is not symmetric under arbitrary rotations, no symmetry forbids such a term, which leads us to expect that it must arise. In all likelihood, it does so through non-Gaussian effects, e.g., a contribution $\theta\langle\theta(\partial_{x}\theta)^{3}\rangle$ from $\cos\theta(\partial_{x}\theta)^{3}$ in (12). Such an average could in principle be nonzero beyond the Gaussian approximation in a polarized state (which would lack $x\to-x$ symmetry).
@@
Our model on a triangular lattice has similar qualitative properties; to see this, however, we have to expand (3) to at least $\mathcal{O}(\nabla^{5})$. Below this, the effect of lattice anisotropy is not evident in the continuum theory. When expanded to $\mathcal{O}(\nabla^{5})$, the term $\propto\mathcal{A}$ doesn’t remain invariant under (11). Further, these have terms that can renormalise the coefficients of $\sin\theta\partial_{y}\theta$ and $\cos\theta\partial_{x}\theta$ differently, making the ordered phases of the model on the triangular and square lattices belong to the same universality class.

### v2

@@ Abstract
We present a detailed analytical and numerical examination, on square and triangular lattices, of the non-reciprocal planar spin model introduced in Dadhichi et al., Phys. Rev. E 101, 052601 (2020). We show that the effect of lattice anisotropy should persist at large scales, leading to a “mass” for the angle field of the spins, and behaviour not in the “Malthusian Toner-Tu” universality class. Numerically, we find evidence of this mass at large values of our non-reciprocity parameter; for smaller values, we find power-law scaling of long-wavelength equal-time correlators in the polar-ordered phase of our lattice model over the system sizes and wavenumber range explored. Focussing on topological defects, we show numerically that defect interactions are highly anisotropic with respect to the mean ordering direction. In particular, the constituents of a ±$\pm 1$ pair are shielded from each other in a class of configurations, deferring their annihilation and allowing time for the nucleation of further defects. The result, we show numerically, is the destruction of the polarised phase via an aster apocalypse reminiscent of that found by Besse et al., Phys. Rev. Lett. 129, 268003 (2022), for the Malthusian Toner-Tu equation.
@@ III The ordered state: Numerics
Figure 2: Statistical properties of the ordered phase on a square lattice. (a) Two-point correlation function $C(r)$ for various system sizes $L$, and $\mathcal{A}=1.5$ with distance $r$ in units of $L/2\pi$, approaches a constant nonzero value for large $r$, indicating long-range order. Inset shows magnetization $m(L)\equiv\sqrt{C(r=L)}$ as a function of $L^{-1}$. (b) Measured structure factors $S_{\perp}$ as functions of $k_{\parallel}$ and $k_{\perp}$ for $L=512$ and $\mathcal{A}=1.5$ show small-$k$ dependences consistent with $k_{\perp}^{-6/5}$ (and $\sim k_{\parallel}^{-2}$ for all but the smallest $k_{\parallel}$). (c) Measured structure factors $S_{\perp}$ as functions of $k_{\parallel}$ and $k_{\perp}$ for $L=512$ and nonreciprocity strength $\mathcal{A}=2.5$ show $k$ dependences consistent with $k_{\perp}^{-6/5}$ except at the smallest $k_{\perp}$ (and $\sim k_{\parallel}^{-2}$ flattening out at the smallest $k_{\parallel}$). The flattening of the structure factor at small $k$ is a signature of the generation of a mass in our model. Since it has to vanish as $\mathcal{A}\to 0$, we expect the mass to be larger for larger $\mathcal{A}$. This expectation is borne out by our observation that the structure factor flattens out at a larger wavenumber as $\mathcal{A}$ is increased. All plots were obtained by averaging over five independent noise realisations and the reduced temperature is $T=0.2$ for each of them. An analogous figure for numerics on a triangular lattice is displayed in Appendix B.
@@
We first examine the orientationally ordered state of our model. Initialising our system in an ordered state, we measure isotropically averaged spatial spin correlations $C(r)=\langle{\bf p}(r)\cdot{\bf p}(0)\rangle_{t}$ the angular brackets with a subscript denotes an average over that variable, for different system sizes $L$, nonreciprocity $\mathcal{A}$, and noise strength $T$. We benchmark our code by measuring these quantities for $\mathcal{A}=0$ which, as expected, displays quasi-long-range order below a defect-unbinding transition at temperature $T_{KT}\simeq 0.8$ for a $128\times 128$ square lattice.. For $\mathcal{A}\neq 0$, we find that $C(r)$ has a non-zero value at large $r$ signalling the existence of long-range order. That, in itself, is not too surprising: as we will discuss in the next section, when the order parameter space and the real space coincide, as it does in our model for $\mathcal{A}\neq 0$, lattice anisotropy explicitly breaks the continuous rotation symmetry of the model to a discrete one, even in equilibrium. Therefore, at low temperatures, we would expect long-range order in any such model. However, the reduction to a discrete symmetry would imply that the static structure factor of angular fluctuations $S_{\perp}({\bf k})\equiv\langle|\hat{\mathbf{p}}_{\perp}({\bf k},t)|^{2}\rangle$, where $\hat{\mathbf{p}}_{\perp}$ is the component of $\hat{\mathbf{p}}$ perpendicular to the instantaneous macroscopic ordering direction and ${\bf k}$ is the wavevector, should go to a $|{\bf k}|-$independent constant at small $|{\bf k}|$. We plot $S_{\perp}(k_{\parallel})$ and $S_{\perp}(k_{\perp})$, where $k_{\parallel}$ and $k_{\perp}$ are the components of ${\bf k}$ along and orthogonal to the ordering direction respectively, in Fig. 2b and c, for two values of $\mathcal{A}$. While for $\mathcal{A}=1.5$, we find a clear departure from the $1/k_{\perp}^{2}$ expected when $\mathcal{A}=0$, they show a dependence on wavenumber consistent with a power law at least in the regime we have explored. The “mass” (see the next section) conferred on the transverse fluctuations by the identification of the order parameter and lattice directions is not detected at these scales at this value of $\mathcal{A}$. In fact, Fig. 2b is consistent with the findings of [20]. In contrast, a distinct flattening of the structure factor at small $k_{\parallel}$ is observed in systems with $\mathcal{A}=2.5$ (to measure this structure factor, we only averaged over realisations in which the ordered phase was not destroyed via an aster explosion—to be discussed below—within the run-time of our simulation). While the flattening is less obvious at small $k_{\perp}$, this is a clear signature of the generation of mass. We discuss the mechanism for this in the next section.
@@ IV From lattice to continuum: the lingering effects of lattice anisotropy
where we have used the fact that the rank $n$ tensor $\sum_{a}[{\bf e}^{a}]^{\otimes n}$ vanishes for odd $n$. Writing $\hat{\mathbf{p}}({\bf x},t)\equiv\{\cos\theta({\bf x},t),\sin\theta({\bf x},t)\}$, where we have defined an angle field $\theta({\bf x},t)$ whose values at lattice sites $i$ correspond to $\theta^{i}$, the equation of motion of [17] on a square lattice to order $\sigma^{3}$ reduces to
@@
Additionally, reflection symmetry ensures that it is invariant under $y\to-y$, $\theta\to-\theta$ [25]. While (10) is symmetric under these transformations for terms up to $\mathcal{O}(\sigma)$, it is clearly not symmetric under (11) for terms at higher orders in $\sigma$. For instance, the term $\propto\sigma^{3}/6$ clearly breaks this symmetry at $\mathcal{O}(\nabla)$ in a way that modifies the scaling behaviour qualitatively (a hyperscaling relation that relates two relevant nonlinearities of the flocking model in isotropic space is no longer valid). Therefore, the microscopic model proposed in [17] cannot describe the spin wave dynamics of flocks in a fully rotation-invariant space. Second, one would expect the coupling of the order parameter to lattice directions to lead to an aligning field that favours alignment along any one of those directions, and thus a restoring torque proportional to the deviation $\theta$ of the orientation from that direction [19]. Such a term is apparently absent in the dynamics as derived above, but fluctuations may generate it though the interaction in [17] doesn’t have a special aligning direction, unlike traditional models with crystal fields [19, 18]. First notice that even though the model is not rotation-invariant, $\theta$ is still a compact, pseudoscalar variable and its equation of motion therefore has to be symmetric under $\theta\to\theta+2\pi$. This implies that the “mass” term, if it exists, must be $\propto\sin\theta$. We notice that the $\mathcal{O}(\sigma^{2})$ term in (10) can be rewritten as $\partial_{x}^{2}\theta+\sin^{2}\theta(\partial_{y}^{2}\theta-\partial_{x}^{2}\theta)$. In a system that does not have rotation invariance and has long-range order, $\sin^{2}\theta(\partial_{y}^{2}\theta-\partial_{x}^{2}\theta)$ can be replaced in a “Hartree” spirit by $\langle\sin\theta(\partial_{y}^{2}\theta-\partial_{x}^{2}\theta)\rangle\sin\theta$. No symmetry dictates that the quantity in angle brackets should vanish, so it should in general be non-zero, imparting a mass to the orientational fluctuations [^6] Note that superficially similar nonlinearities arise in the dynamics of polar and nematic active systems as well; for instance, at $\mathcal{O}(\nabla^{2})$, nonlinearities that remain invariant under the transformation (11) include $\{\sin 2\theta[(\partial_{x}\theta)^{2}-(\partial_{y}\theta)^{2}]-2\cos 2\theta\partial_{x}\theta\partial_{y}\theta\}$ and $[\cos 2\theta(\partial_{x}^{2}\theta-\partial_{y}^{2}\theta)+2\sin 2\theta\partial_{x}\partial_{y}\theta]$ with independent coefficients. Naively, it would seem that a term $\propto\sin 2\theta$ may be generated from $\sin 2\theta[(\partial_{x}\theta)^{2}-(\partial_{y}\theta)^{2}]$ in a state with long range order. It would also seem that another such mass term may be generated from $\cos 2\theta(\partial_{x}^{2}\theta-\partial_{y}^{2}\theta)$. Of course, no such term is generated because unlike both of these nonlinearities, a term $\propto\sin 2\theta$ or $\sin\theta$ is not invariant under (11). In contrast, such terms can be generated in (10) because it is not invariant under (11).. In Appendix B we display the static structure factor $S_{\perp}$ of angular fluctuations obtained numerically using the model of [17], which clearly shows that a mass is generated. This places the models considered in [17, 15, 16] within the active clock model universality class. It was conclusively shown in [18] that, unlike passive clock models which at high enough lattice coordination numbers and at finite noise strengths behave as isotropic spin systems, active $q$-state clock models remain only $q$-fold symmetric, and are thus distinct from polar active systems in isotropic space, for any $q$ no matter how large. Since order in the presence of anchoring to lattice directions should be long-ranged, the quasi-long-range ordered state reported in [15] is likely to be a finite-size effect.
@@
Since the interaction in [17] corresponds to the one we consider here to $\mathcal{O}(\sigma)$, and since the terms we identified as breaking isotropy all arise at higher order in $\sigma$, one might imagine that the spin wave theory of our model belongs to the Malthusian Toner-Tu universality class. We now argue that this is not the case. To do so, we use (3) to obtain the continuum version of our model (7) to $\mathcal{O}(\nabla^{4})$ on square lattices, yielding the equation of motion
$\partial_{t}\theta=2\Bigg{(}\mathcal{A}(\cos\theta\partial_{x}\theta+\sin\theta\partial_{y}\theta)+\frac{J}{2}\nabla^{2}\theta+\frac{\mathcal{A}}{6}\{\cos\theta[\partial_{x}^{3}\theta-(\partial_{x}\theta)^{3}]+\sin\theta[\partial_{y}^{3}\theta-(\partial_{y}\theta)^{3}]\}\\ +\frac{J}{24}\{\partial_{x}^{4}\theta+\partial_{y}^{4}\theta-6[(\partial_{x}\theta)^{2}\partial_{y}^{2}\theta+(\partial_{y}\theta)^{2}\partial_{x}^{2}\theta]\}\Bigg{)}+\sqrt{2T}\xi_{\theta}\,$ (12)
@@
for the angle field. The terms $\propto\mathcal{A}/6$ and $J/24$ in the large parentheses are not symmetric under (11), for distinct reasons. First, we discuss the term $\propto J/24$ which is symmetric under $\theta\to\theta+\theta_{0}$. This is because space and spin are decoupled in our model when $\mathcal{A}=0$, and spin rotation symmetry is preserved irrespective of the spatial anisotropy. The spatial gradient structure reflects the four-fold anisotropy of the lattice. The term $\propto\mathcal{A}/6$ is more interesting: the nonlinearities $\cos\theta(\partial_{x}\theta)^{3}$ and $\sin\theta(\partial_{y}\theta)^{3}$ have structures that can renormalise the coefficients of the terms $\cos\theta(\partial_{x}\theta)$ and $\sin\theta(\partial_{y}\theta)$ implying that the coefficient of the terms $\sin\theta\partial_{y}\theta$ and $\cos\theta\partial_{x}\theta$ need not remain equal in a renormalised theory, breaking the symmetry of our model under (11) even at $\mathcal{O}(\nabla)$. Since both of these terms contain relevant nonlinearities [25] (with a symmetry-mandated hyperscaling relation keeping the ratio of their coefficients fixed under renormalisation in flocks in isotropic space), the hydrodynamic behaviour of the ordered phase of our model cannot belong to the Toner-Tu universality class.
Whether a term $\propto\sin\theta$ is generated in our model is more complicated to resolve analytically. As the dynamics is not symmetric under arbitrary rotations, no symmetry forbids such a term, which leads us to expect that it must arise. In all likelihood, it does so through terms of the form $\sin\theta(\partial_{y}\theta)^{3}$ where the average $\langle(\partial_{y}\theta)^{3}\rangle$ could in principle be nonzero beyond the Gaussian approximation. The fact that our model does have a mass that increases with $\mathcal{A}$ is however borne out by our numerical simulations [^7] After our article was uploaded on arXiv, another preprint examining this class of models and reaching similar conclusions regarding the mass generation and the lingering effects of anisotropy on hydrodynamics, appeared [26].
@@
Our model on a triangular lattice has similar qualitative properties; to see this, however, we have to expand (3) to at least $\mathcal{O}(\nabla^{5})$. Below this, the effect of lattice anisotropy is not evident in the continuum theory. When expanded to $\mathcal{O}(\nabla^{5})$, the term $\propto\mathcal{A}$ doesn’t remain invariant under (11). Further, these have terms that can renormalise the coefficients of $\sin\theta\partial_{y}\theta$ and $\cos\theta\partial_{x}\theta$ differently, making the ordered phases and ordering transitions of the model on the triangular and square lattices belong to the corresponding active clock model universality classes.